OpenStudy (anonymous):

(a + b)2 = a2 + b2 If the statement is false, give a counterexample. ____________________________________________ A) False; (2 + 0)2 = 4 B) False; (1 + 2)2 = 9 C) False; (0 + 1)2 = 1 D) The statement is true.

OpenStudy (anonymous):

Well I know it's false, but I'm pretty sure all the ones with false are right.. ?!

OpenStudy (lgbasallote):

lol do it like the question A. (2 + 0) ^2 = 2^2 + 0^2 = 4 B. (1 + 2)^2 = 1^2 + 2^2 = 9 C. (0 + 1)^2 = 0^2 + 1^2 = 1 remember you are looking for double false

OpenStudy (anonymous):

Double false..?!

OpenStudy (lgbasallote):

your objective is to prove (a+b)^2 = a^2 + b^2 WRONG

OpenStudy (anonymous):

Mhmm.

OpenStudy (lgbasallote):

by double false i meant that \((a+b)^2 \ne a^2 + b^2\)

OpenStudy (lgbasallote):

it's double false because it has a FALSE at the beginning as well...dont mind it...that was supposed to be a pun

OpenStudy (anonymous):

Oh! Okay. But I solved them, I mean 2+0 = 2. 2^2= 4 . So that can't be right. A isn't right. Right?

OpenStudy (lgbasallote):

uhmmm if you mean A isnt the answer then yes

OpenStudy (anonymous):

Okay yeah I meant that. XD. 1+2= 3. 3^2= 9. So its not B either.

OpenStudy (anonymous):

0+1= 1. 1^2 = 1. So its not C. I clicked D but they said it was wrong. So now I'm like. Confused.

OpenStudy (lgbasallote):

i think you are confused...look at what i did (1 + 2)^2 = !^2 + 2^2

OpenStudy (anonymous):

Oh............ Heh. okay nvm...... Lol. XD

OpenStudy (anonymous):

It's B!

OpenStudy (lgbasallote):

you shouldve verified with me before clicking something :P lol but at least you learned something ;D

OpenStudy (anonymous):

Heh yeah. :D